OpenStudy (anonymous):

a ball is thrown vertically upward from the top of a cliff. the height of the ball is modelled by the function h(t) = 65 + 10t - 5t^2, where h(t) is the height in metres and t is the time in seconds. determine when the ball will reach its max height

OpenStudy (anonymous):

To solve for the vertex (peak of arc): -b --- 2a This will give you the time at peak, if you need to know the peak value, plug that answer into x in the original problem and solve for y, or peak elevation of arc.

OpenStudy (anonymous):

@hpilato

OpenStudy (anonymous):

well im supposed to factor the equation.. so i would have -5(-13-5t+t^2)... but im stuck there

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

there's no need to factor use ehuman's method

jimthompson5910 (jim_thompson5910):

use the formula t = -b/(2a) in this case, a = -5 and b = 10 t = -b/(2a) t = -10/(2*(-5)) t = -10/(-10) t = 1 So it will take 1 second for the ball to reach the max height.

OpenStudy (anonymous):

ok! thank you both so much, i wouldnt have known to do that! :)

OpenStudy (anonymous):

it is derived from the quadratic formula, and gives the x value of the vertex.